150g of a gas at 10.0°C and 1.0 atm pressure occupies 913L volume. The volume of the gas increases to 988L and its temperature increases to 41.5°C while the pressure remains constant. Calculate ΔE, q and w for the gas. (Specific heat of the gas = 5.2J/g°C and 1L.atm = 101.3J).
1
Expert's answer
2020-05-25T14:13:18-0400

W = p∆V

W = 101,325 Pa * (988 - 913) x10-3 m3 = 7599.375 J or 7.6 kJ work done by a gas


Q = mcp∆T

Q = 150 g * 5.2J/g°C * (41.5 - 10) = 24570 J or 24.6 kJ


∆E = Q - W = 7.6 -(-24.6) = 32.2 kJ